OpenStudy (anonymous):

how do i convert 1.46e6 lbm/in^2 to lnf/in^2?

OpenStudy (amistre64):

what do lbm and lnf stand for?

OpenStudy (amistre64):

when you know the ratio of one to the other then its just a matter of division

OpenStudy (anonymous):

lbm = like constant weight lbf = weight with force added to it by gravety lbf/in^2 = psi

OpenStudy (amistre64):

if, 1 lbm = K lnf then \[\frac{a~\cancel{lbm}}{in^2}*\frac{K~lnf}{\cancel{lbm}}=\frac{aK~lnf}{in^2}\]

OpenStudy (amistre64):

constant weight, you mean mass?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

i feel like im coming in at the tail end of a problem .... what was the original setup?

OpenStudy (anonymous):

A simply supported basswood beam is 30 inches long and carries a load of 210 pounds. It has a hollow box cross-section where b = 5 inches, h = 8 inches, and t = 0.5 inches. Basswood has the following material properties: Modulus of Elasticity = 1.46x106lbm/in2 and Density = 28lbm/ft3. Determine the moment of inertia, deflection, volume, and weight of the beam. moment of intertia is found to be 99

OpenStudy (amistre64):

modulus of elasticity has to deal with stress; which is Force/(Cross sectional Area) right?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

and density is mass per volume

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

would i just multiply by 32.2?

OpenStudy (amistre64):

well, Force = mass x acceleration; so mass X g would be my idea as well

OpenStudy (amistre64):

28lb = mass x 32.2 mass = 28/32.2

OpenStudy (amistre64):

lb is a force already, so to define mass we would have to divide out the acceleration

OpenStudy (anonymous):

when i use the equation to solve for deflection, i am not able to get the correct answer using the elasticity value

OpenStudy (amistre64):

what is the deflection formula? it been awhile since physics :)

OpenStudy (anonymous):

(PL^3)/(48EI)

OpenStudy (amistre64):

lol, this doesnt really help work out the theory, but it seems to be a calculator for deflection. im out of time at the library so i gotta head out http://www.botlanta.org/converters/dale-calc/bending.html

OpenStudy (anonymous):

alright thanks for your help